The energy required to break a double C=C covalent bond is about 1.13x10^-18J. Calculate the wavelength of the photons that can supply the specified energy. It is well to note that the Planck's constant is 6.6256x10^-34J/s and the speed of light equals 3x10^8m/sec. The Avogadro's number is 6.022x10^23 per mole.
Answer #1

Solution:

E = h * c / ?

=> 1.13 * 10^-18 = 6.625 * 10^-34 * 3 * 10^8 / ?

=> ? = 17.588 * 10^-8 m

=> ?= 1.758 nm

please comment for queries.

Please rate thanks.